A boat travels 60 km upstream (against the current) in 5 hours. the boat travels the same distance downstream in 3 hours. What is the rate of the boat in still water? What is the rate of the current?

Answer by mananth(16946)

let rate of boat in still water = x
let rate of current be y
..
Speed = distance / time
..
speed upstream = x-y
speed downstream = x+y
..
time upstream = 60/x-y =5
60/5 = x-y
x-y =12...............1
time downstream = 60/x+y = 3
x+y =20...............2
add equation 1 & 2
2x = 12+20
2x = 32
x=16 mph rate of boat.
x-y =12
so y=4 rate of current